An experiment was conducted to measure the weight of the water in a cylindrical can. The
water drained out through a small hole in the bottom of the can. Find the quadratic best
fit to the following data obtained from the experiment:

Time (in seconds) Weight of 6 11 16 21 26 31 36 41 47 water in 2.85 2.74 2.63 2.46 2.35 2.24 2.14 2.03 1.97 1.92 Newtons)
Answer #1

The general foim o quadratis best fit is given by ga at batch² ->0 where, a,b and care constants to be determined The NormalG substitute above values in 0.3 . weget 외 coa + 236 b + 7678C = 23.33 → Czrn=10) 236 9 + 7678 b + 2816720 = 505.60 (6 76789